### v2

# A foundation machine learning potential with polarizable long-range interactions for materials modelling

Abstract
Long-range interactions are essential determinants of chemical system behaviour across diverse environments. We present a novel framework that integrates explicit polarizable long-range physics with an equivariant graph neural network potential. It employs a physically motivated polarizable charge equilibration scheme that directly optimizes electrostatic interaction energies rather than partial charges. The foundation model, trained across the periodic table up to Pu, demonstrates strong performance across key materials modelling challenges. It effectively captures long-range interactions that are challenging for traditional message-passing mechanisms and accurately reproduces polarization effects under external electric fields. We have applied the model to mechanical properties, ionic diffusivity in solid-state electrolytes, ferroelectric phase transitions, and reactive dynamics at electrode-electrolyte interfaces, highlighting the model’s capacity to balance accuracy and computational efficiency. Furthermore, we show that as a foundation model, it can be efficiently finetuned to achieve high-level accuracy for specific challenging systems.
Introduction
Molecular dynamics (MD) simulations are indispensable in describing different phenomena at the atomic level in disciplines such as chemistry, materials science, and biology 1, 2, 3, 4. While ab initio MD (AIMD)5 simulation approaches offer unparalleled accuracy, their computational demands constrain their application to small timescales and length scales. On the other hand, classical MD (CMD)6, 7 simulations offer computational efficiency, but their accuracy is contingent upon empirical parameters. Machine learning interatomic potentials (MLIPs)8, 9, 10, 11 provide a solution that balances accuracy and computational efficiency, outperforming traditional methods. In particular, MLIPs based on equivariant graph neural networks, such as NequIP12, DimeNet13, and MACE14, achieve excellent performance by introducing equivariant and invariant symmetries. Moreover, universal models trained on the periodic table such as M3GNet15, CHGNet16, and GNoME17 have emerged, showing remarkable prospects for materials discovery.
While existing local MLIPs with a cut-off of around 5 Å perform well in simulating interactions within localized chemical environments, they may fail to capture long-range phenomena. This hinders their ability to understand and elucidate the behaviours of complex materials 18. An approach is to implicitly incorporate the effects of long-range interactions through message-passing mechanisms 19, which can extend the cutoff through layer-wise propagation. However, if certain parts of the system are disconnected on the graph, such as two molecules separated by a distance beyond the cutoff value, the message-passing scheme becomes ineffective 20. A promising direction involves investigating the potential advantages of explicitly including long-range interactions in the model formulation. These behaviours encompass electrostatic interactions, involving forces between charged particles, and dispersion terms, which arise from dynamic fluctuations in electron distribution within molecules or atoms.
Various explicit models have been proposed to address the challenges associated with long-range interactions by including electrostatic interactions. In these methods, charges are usually obtained through charge equilibration (QEq) 21 principles, including the Charge Equilibration via Neural Network Technique (CENT) method 22, Fourth-generation High-Dimensional Neural Network Potential (4G-HDNNP) 23, 24. Generally, electrostatic parameters were trained to minimise the difference between QEq charges and reference partial charges derived from quantum mechanics (QM) calculations. Notably, partial charges obtained from QM calculations are either derived from the direct partitioning of the molecular wave function into atomic contributions or calculated based on the analysis of physical observables. However, due to the incompleteness of basis sets and variances in the partitioning methods, properties calculated based on partial charges may not always be reliable 25. The ambiguity in DFT-assigned charges suggests that directly learning of them may be inessential for constructing accurate interatomic potentials. On the contrary, the polarizable charge equilibration (PQEq) 25, 26, 27 method proposed by Naserifar et al. enhances the QEq method by using QM electrostatic interaction energies instead of partial charges as targets for evaluating interatomic potentials. Additionally, PQEq has proven to be effective in capturing polarization effects by explicitly introducing core-shell displacement. Another way to solve the partition-dependent issue is to take dipole moment into the loss function, as done by SpookyNet 28. However, it still uses point-like charges, making it behaves like QEq-based methods when dealing with polarization effects.
In this work, we introduce a novel framework that integrates the equivariant message-passing neural network potentials with the explicit polarizable long-range potential. We initially conducted benchmark tests on datasets containing systems with various charge states proposed by references 24, 29, showcasing the efficacy of our framework in the handling both long-range interactions and varying charge environments. Building on these results, we extended the framework to develop a foundation model encompassing the periodic table up to Pu, while maintaining computational efficiency. We have applied this foundation model across different areas of applications, including the predictions of mechanical properties of materials, phase transitions in ferroelectric materials, and reactive molecular dynamics of solid-state batteries interface, thus leading to significant advancements in the fields of materials science and chemical simulations. The model also demonstrates transferability, accurately predicting interactions between clusters in various charge states, and molecular polarizability. Furthermore, we show that the foundation model can be efficiently finetuned to achieve ab initio accuracy for specific challenging systems.
Results
Theoretical framework
Figure 1: a, Overview of the framework. The framework takes atomic numbers ($Z_{i}$) and coordinates ($\bm{r}_{i}$) as inputs and integrates neural network block and explicit polarizable long-range interactions block. For the neural network block, atomic numbers are converted into feature vectors via one-hot encoding, while the coordinates are transformed into a representation of the local environment using radial basis functions and spherical harmonics to capture the distance ($|\bm{R}_{ij}|$) and directional ($\bm{R}_{ij}/|\bm{R}_{ij}|$) information. These embeddings are processed through an equivariant graph neural network to output the scalar MLIP energy $E_{0}$. Explicit, physically-grounded calculations are performed for long-range effects. An electrostatic energy component ($E_{\rm PQEq}$) is calculated using the polarizable charge equilibration (PQEq) method, and a dispersion energy component ($E_{\rm D3}$) is calculated using the DFT-D3 method. The total energy is the sum of these components. Gradients of the energy yield forces and stress on the system. The long-range block also provides partial charges as an auxiliary output. b, A flowchart detailing the iterative self-consistent procedure for the PQEq method. Starting with atomic core positions ($\bm{r}_{ic}$), atomic numbers ($Z_{i}$), and predetermined atomic parameters (electronegativities $\chi_{i}^{0}$, chemical hardness $\eta_{i}^{0}$, and spring constant $K_{s}^{i}$), the method uses a core-shell Gaussian model to build and solve a system of linear equations. This loop is repeated until the partial charges ($q_{i}$) and shell positions ($\bm{r}_{is}$) converge, yielding the final electrostatic energy ($E_{\rm PQEq}$) and the converged partial charges. c, Partition of core-shell Gaussian charge model used in long-range electrostatics. The partial charge ($q_{i}$) of each atom is represented as sum of the core ($q_{ic}$) and shell ($q_{is}$) charge. The harmonic spring constant, $K_{s}^{i}$, couples the core and shell.
Our framework, which integrates neural network potential and polarizable long-range interactions, is depicted in Fig. 1. For a given chemical system and boundary conditions, our goal is to construct a mapping from atomic coordinates $\bm{r}_{i}$ and atomic types $Z_{i}$ to the total potential energy $E_{\rm pot}$. The potential energy is expressed as the sum of the second-order expansion 25 with respect to charge fluctuations and the density functional theory 30 D3 (DFT-D3) van der Waals dispersion energies correction 31 term $E_{\rm D3}$,
$\begin{split}E_{\rm pot}&=\sum_{i}\left(E_{0}^{i}(\bm{r}_{i},z_{i})+\chi_{i}^{0}q_{i}+\frac{1}{2}\eta_{i}^{0}q_{i}^{2}+\frac{1}{2}K_{s}^{i}\bm{r}_{ic,is}^{2}\right)\\ &+\sum_{i>j}C_{ik,jl}(\bm{r}_{ik,jl})q_{ik}q_{jl}+E_{\rm D3}\\ &=\sum_{i}E_{0}^{i}+E_{\rm PQEq}+E_{\rm D3}.\end{split}$ (1)
The zeroth-order atomic energy $E_{0}^{i}$ corresponds to the final layer scalar output of the equivariant graph neural networks, while the higher-order self- and interatomic Coulomb interactions represent the polarizable long-range electrostatic interactions $E_{\rm PQEq}$. To account for charge transfer and polarization effects, partial atomic charge $q_{i}$ of atom $i$ is the sum of the nuclear core charge $q_{ic}$ and shell charge $q_{is}$, both of which assume a Gaussian charge distribution form. The first-order coefficients $\chi^{0}$ are electronegativities, commonly defined as half of the sum of ionization potential (IP) and electron affinities (EA). The second-order coefficients $\eta^{0}$ signifies idempotential or chemical hardness, defined as IP-EA. The spring constant $K_{s}^{i}$ denotes the isotropic harmonic connectivity between the shell position $r_{is}$ and core position $r_{ic}$ (equal to $r_{i}$) of atom $i$. The Gaussian electrostatic energy is given by $C(\bm{r}_{ik,jl})q_{ik}q_{jl}$, where $i$ and $j$ are the atomic indices, and $k$ and $l$ represent the core ($c$) or shell ($s$), respectively. The displacement vector is given by $\bm{r}_{ik,jl}=\bm{r}_{ik}-\bm{r}_{jl}$. The detailed derivation of the response to the external electric field can be found in the Supplementary Information. In previous works, PQEq parameters for 102 elements have been derived from experimental data or high-level QM calculations 25, and have accurately reproduced QM electrostatic interaction energies. In the equivariant neural networks 12, the initial features are generated using a trainable one-hot embedding that operates on the atomic types. The interatomic distance of atom $i$ and atom $j$, denoted as $|\bm{R}_{ij}|$, is expanded by radial basis functions. Concurrently, the directional component of the interatomic vectors, expressed as $\bm{R}_{ij}/|\bm{R}_{ij}|$, is expanded by spherical harmonics functions. These features are utilized to construct the atomic graph. Then, through the equivariant message passing 19 schemes, the atomic features are updated. A multilayer perceptron layer is then connected to derive the scalar output. It should be noted that the neural network potential component is specifically designed to capture energy contributions distinct from electrostatic interactions, thereby ensuring no overlap in energy accounting between different components. Subject to the conservation of the net charge and the equality of chemical potentials for all atoms, the linear equations can be used to updated partial charges ($q_{i}$) and shell positions ($\bm{r}_{is}$). The partial charges can be obtained self-consistently in each MD step. The forces on atoms and virial stress on the cell can be calculated via automatic differentiation after partial charges are determined as detailed in Supplementary Information.
Validation on diverse charge-state datasets
[…]
Table 1: The comparison of root means square test error metrics on the different charge-state datasets.
Dataset 4G-HDNNP Maruf’s NequIP This work
$\rm{C_{10}H_{2}/C_{10}H_{3}^{+}}$ Energy (meV/atom) 1.194 1.33 0.44
Force (eV/Å) 0.078 0.071 0.023
[…]
Force (eV/Å) 0.033 N.A. 0.005
[…]
Force (eV/Å) 0.032 2.145 0.028
Au2-MgO Energy (meV/atom) 0.22 1.03 0.14
Force (eV/Å) 0.066 0.034 0.021
BTA-Cu Energy (meV/atom) N.A. 0.48 0.30
Force (eV/Å) N.A. 0.008 0.004
BTA (H2O)-Cu Energy (meV/atom) N.A. 0.71 0.19
Force (eV/Å) N.A. 0.010 0.005
For the carbon chains, the introduction of additional H+ causes some C atoms to exhibit opposite charge states, as shown in Fig. S1 a. Although we did not fit partial charges, our model qualitatively captures the distribution of DFT partial charges, as shown also in Fig. S1 a. In the linear C10H2 molecule, our model demonstrates symmetric charge distribution around the centre which shows agreement with DFT. After explicitly incorporating physical long-range interactions, our framework achieved additional improvements in error metrics compared to NequIP and 4G-HDNNP. For positively charged NaCl clusters, the predicted energy and forces by our model are still more accurate compared to 4G-HDNNP, as well as the potential energy surface shown in Fig. S1 b. For Au2 adsorption on MgO surfaces and benzotriazole interactions with Cu-(111) surfaces in dry and aqueous environments, our model outperforms NequIP significantly. Additionally, our model correctly predicted the preferential adsorption geometries for both doped and undoped MgO, as listed in Table S1. As for the charged Ag clusters, due to their small system size and lacking long-range effects, the entire structure falls within the typical cutoff radius used by the MLIP. The energy dependence on the overall charge state of clusters leads to degeneracy between atomic structures and potential energy surfaces, resulting in poor performance of charge-independent NequIP model. We acknowledge that for such small, highly charged systems, 4G-HDNNP approach to learning charge equilibration parameters is more convenient. Nonetheless, we also achieved reduced force error with the fixed physical parameters compared to 4G-HDNNP. Although PQEq partial charges are not directly comparable to DFT Hirshfeld ones, we compare them to show their relevance. Varying degrees of agreement across different chemical systems are observed as shown in Table S2. Our comprehensive evaluation demonstrates that explicit incorporation of physical long-range interactions significantly enhances the performance of MLIPs across diverse charge-state systems. Notably, our proposed framework outperforms 4G-HDNNP in most cases for energies and forces without fitting partial charges. This advantage likely stems from a more physically meaningful partition of the total energy.
Foundation model benchmark
Figure 2: a, The interaction energies for Na-Na (green) and Na-Cl (purple) dimers predicted by our model, w/o-lr model, and reference calculations using DFT. b, Evaluation of polarizable interactions in water molecules. The water molecule model oriented in the yz-plane, with oxygen and hydrogen atoms shown in red and white, respectively. The energy response curves as a function of external electric field ($\bm{\epsilon}$) strength applied along the x-axis. Results are compared among our model with polarizable long-range interactions (blue), the QEq-based model (yellow), and the reference DFT calculations (green) serving as the reference values. c-e, Comparison of bulk modulus from first principles calculations and our model: c, Voigt approach, d, Reuss method, and c, Hill average.
We trained a foundation model for all the periodic table elements up to Pu using MPtrj dataset 16, as described in the Methods section, to evaluate the feasibility of this framework. We also trained a model maintaining the same MLIPs framework and dataset but without the polarizable long-range interactions (w/o-lr model). The former one demonstrated mean absolute errors (MAE) of 18 meV per atom for energies, 0.065 eV/Å for forces, and 0.301 GPa for stresses across the test set, outperforming the w/o-lr model, as shown in Table S4. These proved the benefit of incorporating long-range interactions in the enhancement of prediction accuracy.
Message-passing mechanisms will fail when graph connections break in the system. For instance, in dimers, when the interatomic distance exceeds the graph cut-off, the atoms no longer interact with each other. We analysed interaction energies for Na-Na and Na-Cl dimers across interatomic distances ranging from 1.8 Å to 12.0 Å. As shown in Fig. 2 a, both models (our model and w/o-lr model) trained on MPtrj dataset accurately reproduce the equilibrium bond lengths of the two dimers. However, the w/o-lr model failed to differentiate interaction energies beyond the MLIPs cutoff distance, i.e., 5.0 Å, while our model successfully captured the DFT potential energy surface throughout the entire distance range, demonstrating its capability in describing extended ionic interactions. Moreover, the message-passing mechanism exhibits deficiencies when simulating layered materials. Taking lithium iron phosphate (LiFePO4) as a case study, we observed that the w/o-lr model predicted an irreversible phase transition as shown in Fig. S3. In contrast, our model correctly reproduced the experimentally 32 and AIMD 33 observed thermally stable olivine structure. These results collectively demonstrate that the explicit long-range interactions are crucial for accurately describing both ionic systems and layered materials, particularly in predicting structural stability and thermal behaviour.
Polarization plays a fundamental role in determining molecular responses to external electric fields. Conventional QEq-based models have limitations in their response to an external electric field applied orthogonally to the molecular dipole. We validated this capability through a benchmark study examining a water molecule’s response to external electrostatic fields 34. The water molecule was fixed in the $yz$-plane, and we applied varying external electric fields along the $x$-axis to derive energy curves, with the results presented in Fig. 2 b. The model with polarizable long-range interactions demonstrates a trend consistent with the reference values obtained from DFT. Due to the external electrostatic field direction consistently being orthogonal to the molecule, the electrostatic energy in the QEq-based model (like 4G-HDNNP) is 0 and is completely independent of field strength ($i.e.,\bm{\mu}\cdot\bm{\epsilon}=0$, where $\bm{\mu}$, $\bm{\epsilon}$ are dipole moment and electric field, respectively). This is because the QEq model treats atoms in the molecule as point charges or Gaussian charges, without distinguishing between core and shell components. We also compared the response of O atom partial charges across DFT Hirshfeld calculations, our model, and the conventional QEq model as shown in Fig. S4. The agreement between PQEq model and DFT is observed, while QEq model completely lacks response. This comparative analysis establishes that the explicit polarizable interactions are crucial for modelling molecular systems subject to external electric fields.
Besides the failure of response in static conditions, the QEq method may exhibit non-physical behaviour in dynamics 35. As illustrated in Fig. S5, we conducted additional investigations of two water molecules under an external electric field of 0.25 V/Å. In charge distribution models, it is essential to allow systems to form an internal electric field opposing the external field. The QEq model assumes the system behaves as a perfect conductor without penalizing charge transfer as a function of distance. This limitation of the QEq method leads to non-physical charge transfer between molecules separated by large distances 35. In contrast, within PQEq model, the individual water molecules can polarize, generating internal electric fields that counteract the external field, thereby capturing more realistic electrostatic responses in molecular systems. Under QEq scheme, the water molecules accumulate non-zero net charges and migrate to the top and bottom of the simulation box. In contrast, our model maintains charge neutrality of the water molecules, which remain stationary despite the applied electric field, aligning with DFT results.
We conducted benchmarking using properties that were not labeled during the training process. First, we applied the foundation model to predict the mechanical properties of 10,154 materials from the Materials Project36. Fig. 2 c-e illustrate the comparison of the bulk modulus $B$ determined by the foundation model and DFT. Our foundation model demonstrated impressive performance, achieving an $R^{2}$ of 0.94 with the Voigt approach37 $B_{\rm V}$ and Hill average method38 $B_{\rm VRH}$, and 0.93 using the Reuss method39 $B_{\rm R}$. In comparison, as shown in Fig. S2, the w/o-lr model achieved lower $R^{2}$ values of 0.89, 0.86, and 0.88 for $B_{\rm V}$, $B_{\rm R}$, and $B_{\rm VRH}$, respectively. These not only highlight the robust performance of the model but also establish a solid foundation for the high-throughput screening of materials with exceptional mechanical properties.
Transferability of the foundation model
In terms of foundation model transferability, evaluations were conducted across diverse systems. The model accurately reproduces the potential energy surfaces and successfully distinguishes between neutral and ionic states in OH-OH systems, capturing the long-range Coulombic repulsion of ionic state as shown in Fig. S8. Further validation using a periodic water system demonstrates the model’s capability to predict polarization effects under varying electric fields, achieving comparable performance to specialized models as depicted in Fig. S9. Additionally, as shown in Fig. S10, when tested on 7,211 molecules from the QM-7b dataset 40, the model shows remarkable accuracy in predicting molecular polarizability with a mean absolute error of 4.57 a.u., highlighting its robust transferability in capturing polarization.
Computational efficiency
Despite the incorporation of additional long-range interaction calculations, our foundation model maintains high computational efficiency. Performance benchmarks were conducted on a single NVIDIA H100 GPU to quantitatively assess the computational cost. As illustrated in Fig. S6, for a periodic system comprising 2160 atoms, our model requires approximately 0.07 s per molecular dynamics time step, representing a significant improvement over conventional universal MLIPs which require approximately 0.91 s. This computational advantage extends to larger systems, as demonstrated by simulations of a 24000-atom system, where our model maintains efficiency with only 1.49 s per time step. Such computational performance enables nanosecond-scale molecular dynamics simulations of systems containing tens of thousands of atoms, making it practical for more realistic modelling of materials.
Materials modelling applications
Figure 3: a, Crystal structure of $\rm{Ia\bar{3}d}$-Li7La3Zr2O12 and Arrhenius plots depicting the lithium-ion diffusion coefficients across varying temperatures. The dark blue polyhedron signifies La located at the 24(c) site and the light brown polyhedron indicates Zr at the 16(a) site. Li fraction occupies the 24(d) and 96(h) sites. Predicted diffusion coefficients of AIMD 41, our model, and w/o-lr model complete with error bars, are presented to calculate activation energies. b, 2-ns mean square displacements using our model of lithium-ion in c-LLZO with different temperatures ranging from 800 K to 1800 K in an increment of 200K. The linear dashed grey lines, with a slope of 1, are also plotted.
To demonstrate the capability of our model to reproduce results from AIMD simulations, we simulate the kinetic transport properties of a solid-state electrolyte. We investigated lithium-ion diffusivity within the cubic phase of $\rm Li_{7}La_{3}Zr_{2}O_{12}$ (c-LLZO)42, a crystalline superionic conductor known for its remarkable stability as a lithium conductor. We conduct MD simulations on the c-LLZO comprising 64 formula units, with a duration of 2 ns for each temperature range from 800 K to 1800K. Fig. 3 a presents the Arrhenius plots derived from our model and AIMD41 simulations. The diffusion coefficients were determined from the slope of the logarithmic mean square displacements (MSD) versus logarithmic time within the Fickian regime43, as depicted in Fig. 3 b. Compared to previous work using AIMD simulations, our model clearly reproduces the diffusion coefficients and activation energy. The performance surpasses that of the w/o-lr model which lacks explicit long-range interactions. With the ability to reach larger model sizes and longer simulation times, our model achieves these results with significantly reduced uncertainties. This capability not only delivers comparable results to AIMD simulations but also enables larger-scale simulations that cannot be assessed by AIMD, enhancing the statistical significance of MD studies. Consequently, the framework developed in this work provides more reliable and comprehensive data for diffusion analysis in solid-state electrolytes.
To verify the ability of our model to handle polarized systems, we performed MD simulations on a typical perovskite ferroelectric material $\rm BaTiO_{3}$. As temperature increases, $\rm BaTiO_{3}$ undergoes a non-trivial phase transition process44. Below 183 K, it adopts a rhombohedral structure with local polarization along the [111] direction. As the temperature rises, it transitions to an orthorhombic phase with polarization along the [110] direction, then to a tetragonal phase with polarization along [100] direction at 278 K. Finally, at 403 K, it converts to a cubic paraelectric phase. Accurately simulating phase changes in ferroelectric substances requires precise potential energy functions that can respond to the small atomic shifts and structural changes, as well as account for the free energy surfaces under the conditions of finite-temperature thermodynamics45. The rhombohedral-orthorhombic-tetragonal-cubic phase sequence of $\rm BaTiO_{3}$ has been extensively studied and reproduced using effective Hamiltonians46, 47, CMD48, 49, and specialized MLIPs models445, 50. To detect subtle lattice distortions and variations in free energy that distinguish different phases, a 10×10×10 supercell of $\rm BaTiO_{3}$ was simulated. This allowed for a detailed analysis of the structural changes that occurred between different phases, providing valuable insights into the properties of materials. The use of larger simulation cells (10×10×10 in our work versus 4×4×4 in Gigli et al. 45) may provide some advantages for phase transition studies. Larger supercells can effectively reduce the impact of temperature fluctuations, enabling better temperature control for sampling the free energy landscape, thereby yielding results with greater statistical significance. Fig. 4 a and b illustrate the simulation results, which clearly identify four distinct phases and the three first-order phase transitions within the simulated temperature range. Below 145 K, the overall average polarization of the supercell aligns with the [111] direction, indicative of the rhombohedral phase. At 145 K, a phase transition to orthorhombic is suggested as the y-component of polarization approaches zero. With further increase to 205 K, the polarization predominantly orients along the x-direction, indicating the tetragonal phase. This phase persists until 285 K, where the cubic paraelectric phase is observed. The obtained phase transition temperatures align closely with those predicted by various models and experiments, as shown in Table 2. Notably, utilizing first principles methods or fitting potential energy surfaces to first principles data often underestimates these temperatures, which is attributed to the approximated exchange-correlation functional48 employed in DFT. Although the predicted phase transition temperatures are lower than the experimentally observed values, the model effectively captures the sequence of phase transitions as measured in experiments, demonstrating the efficacy of the foundation model in modeling highly polarized phase transitions in ferroelectric materials.
Figure 4: The temperature dependence of a, lattice constants, and b, local polarizations of unit cells exhibit notable changes during the phase transitions observed from MD simulations on 10×10×10 supercell of BaTiO3.
Table 2: Comparison of BaTiO3 phase transition temperatures obtained by different methods and experiments.
Method $T_{c,\rm{R-O}}$ (K) $T_{c,\rm{O-T}}$ (K) $T_{c,\rm{T-C}}$ (K)
This work (10×10×10) 145 205 285
Effective Hamiltonian (12×12×12) 51 150±10 195±5 265±5
Effective Hamiltonian (16×16×16) 47 119 158 257
Second Principles (16×16×16) 52 140 180 224
ReaxFF (6×6×6) 53 N.A. N.A. 240
MLIPs (4×4×4) 45 18.6 91.4 182.4
Experiments54 183 278 403
The final application studied in this work involves all-solid-state batteries, which represent a breakthrough in the evolution of next-generation energy solutions, owing to their superior energy density and inherent safety features55. Lithium thiophosphate, known for its exceptional ionic conductivity ($\sim 10^{-3}$ S/cm), is deemed the most promising candidate for solid electrolytes and has been widely studied through experiments56, 57, 58 and computer modelings59, 60, 61. Notably, the nanoporous $\beta-\rm{Li_{3}PS_{4}}$ has been validated to exhibit outstanding cycling stability62, presumably attributed to the formation of Li2S and Li3P solid-electrolyte interphases (SEI) during the initial battery cycles, which serves to passivate further degradation of the electrolyte60, 63, 64, 65. To investigate the formation of SEI, we utilized the foundation model to simulate the interfacial MD of the solid electrolyte $\beta-\rm{Li_{3}PS_{4}}$ in contact with the lithium metal anode, encompassing a system of 13,760 atoms.
Figure 5: a, The initial structure of $\beta$-Li3PS4 [010]/Li [001] interface. The elements are colour-coded: Li in green, S in purple, and P in pink. A segment of the primary reaction zone, ranging from 8 to 26 nanometres is displayed. b, Snapshots at different times during the MD simulations. The partial charges on lithium ions are represented with colour coding to enhance the visibility of structural transformations during the formation of SEI. The SEI layer, approximately 8.5 nm in thickness, forms after 4 ns MD simulations, comprising an amorphous $\beta$-Li3PS4/Li2S interface ( 2 nm), a crystalline Li2S layer ( 4.5 nm), and an amorphous Li2S (Li3P)/Li interface layer ( 2 nm) in sequence. c, The distributions of the Li partial atomic charges along the x-direction at initial, 20, 100, 2000, and 4000 ps state. d, The radial distribution function plots of Li-S and Li-P within the electrolyte and SEI layer.
Figure 6: The snapshots capture the dynamic decomposition behavior of PS4, showcasing: a, the sulfur, and b, the phosphorus component. To elucidate the formation mechanism of the SEI layer , atomic component analyses of the c, crystalline, and d, amorphous Li2S regions are compared among 4-ns (upper), 2-ns (middle), and the initial state (lower). The elements are color-coded: Li in green, S in purple, and P in pink. The x-component root mean squared displacements (RMSDx) of Li, P, and S between the initial and the 2-ns states and between 4-ns and 2-ns states are also shown.
The initial $\beta-\rm{Li_{3}PS_{4}}$/Li interfacial structure, measuring 33.15 nm × 3.1 nm × 2.6 nm, was utilized in the MD simulations, as depicted in Fig. 5 a. The highly reactive Li metal began to interact with PS4 tetrahedron at the interface, triggering the formation of the SEI layer as illustrated in Fig. 5 b. With the growth of the SEI layer, electrons were transferred from the Li anode to the $\rm{Li_{3}PS_{4}}$ electrolyte, which causes a gradual increase of the partial charges of Li as they migrate from the anode, across the SEI, and into the electrolyte. This results in a transition from metal lithium to ions. The dynamic behavior of partial charges at the anode-electrolyte interface is a critical aspect showcased in Fig. 5 c. Initially, at 10 ps, the partial charges of lithium changed almost linearly from the anode to the electrolyte. Over time, this distribution evolved and showed a distinct plateau at 100 ps. This phenomenon could be attributed to the ordered structuring of the SEI, signifying the nucleation of crystalline structures. Ultimately, a stable lithium partial charges plateau was formed between the anode and the electrolyte, spanning a range of 15 to 21.5 nm. This is further supported by the visualized structure depicted in Fig. 5 b, where the plateau within the SEI aligns with the Li2S crystal region. Chen et al.59 employed AIMD simulations to investigate the radial distribution functions (RDF) of Li-P and Li-S bonds, which concluded that the decompositions of electrolytes in lithium thiophosphates are primarily due to the decomposition of PS4 tetrahedron by the active lithium metal, leading to the formation of new Li-P and Li-S bonds. Our RDF analysis confirmed the formation of Li-S and Li-P bonds within the SEI and electrolyte as depicted in Fig. 5 d. In agreement with the AIMD simulations59, we observed the emergence of a new Li-P peak within the SEI at approximately 2.5 Å. Furthermore, within the SEI, the Li-S bond displayed RDF characteristics consistent with crystalline Li2S66. For the Li-P bonds, the peak shape closely resembles that of amorphous Li3P AIMD simulations67. As illustrated in Fig. 6 a and b, the ultimate decomposition products of P form only short-range ordered structures, in contrast to S, which forms long-range ordered structures. This observation agrees with the measurements where Li2S crystals were detected but Li3P crystals were absent in cryogenic transmission electron microscopy (cryo-TEM) experiments, as noted in reference57. To enhance the understanding of the formation mechanism, the atomic compositions of the SEI in both crystalline and amorphous Li2S regions were examined.
As depicted in Fig. 6 c and d, within the first 2 ns, the emergence of amorphous Li2S and Li3P regions was attributed to the swift diffusion of P and S atoms. Conversely, the formation of crystalline Li2S regions was predominantly influenced by the swift diffusion of Li atoms and slower diffusion of P and S atoms. As the system further evolves to 4 ns, the diffusions slow down, which also confirmed that the formation of crystalline Li2S region hindered further atomic diffusion and thus slowed the growth of the SEI. As illustrated in Fig. 5 b, during the initial phase of decomposition (10-100 ps), the swift diffusion of P, S, and Li led to the formation of an interface of about 3 nm thick. However, in a later stage between 100 and 2000 ps, the formation of Li2S nuclei impeded the diffusion of P and S atoms, resulting in a period of sluggish SEI growth. Consequently, the interface experienced only a slight expansion in this period, increasing by only 1 nm. Furthermore, in the following 2 ns, there was essentially no SEI growth. The final interface structure ($\sim 8.5$ nm) could be characterized by a 4.5 nm crystalline Li2S region sandwiched between two 2-nm transitional layers. The cessation of growth at this interface suggests that the crystalline Li2S and amorphous Li3P within the crystal region contribute to the stabilization of both the lithium metal anode and the electrolyte, which agrees with the theoretical predictions made by DFT calculations63.
Additionally, we studied the interface reactions between $\rm Li_{6}PS_{5}Cl$ and lithium metal, as detailed in Supplementary Information. As shown in Fig. S12, at a pressure of 1 atm and temperature of 300 K, $\rm Li_{6}PS_{5}Cl$ formed an interface of about 5-nm thick after 2 ns. This indicates that the $\rm Li_{6}PS_{5}Cl$ electrolyte exhibits a higher stability against lithium metal at room temperature compared to $\rm Li_{3}PS_{4}$. Upon elevating the temperature to 500 K, we detected Li2S at the grain boundaries, which aligns with the cryo-TEM observations 57.
Finetuning for enhanced accuracy
[…]
Discussion
In this study, we introduced a novel framework that integrates equivariant machine learning interatomic potentials with explicit long-range polarizable interactions. QEq-based models inherently struggle with atomic polarization because they do not distinguish between core and shell components. Due to their inability to handle atomic-level responses to external fields, they may exhibit non-physical behaviours in some cases as described in the reference 35. Our framework incorporates polarizable long-range interactions and accurately captures atomic polarization in the presence of external electric fields. Since our framework decomposes the total energy into electrostatic energy and remaining energies that only depend on atomic types and positions, any future machine learning techniques and/or high-order electrostatic energy expansion schemes can be readily integrated into it, empowering possible accuracy enhancement.
Our foundation model successfully reproduced DFT potential energy surfaces for Na-Cl/Na-Na dimers across extended interatomic distances, outperforming conventional message-passing MLIP beyond the cut-off regions. The model demonstrates exceptional versatility across multiple materials science challenges, including accurate prediction of thermal stability in complex layered materials, high-fidelity capture of molecular responses to external electric fields, and reliable prediction of bulk modulus. The model has been also successfully applied to study dynamic properties and reactive molecular dynamics, such as lithium-ion diffusion in solid-state electrolytes and temperature-dependent phase transitions in ferroelectric materials like BaTiO3, while also elucidating interface formation mechanisms between lithium thiophosphate electrolytes and lithium metal anodes in solid-state batteries.
Beyond these specific applications, we evaluated the model’s transferability, a critical attribute for a foundation model. The model accurately reproduces the potential energy surfaces for OH-OH systems, correctly distinguishing between the neutral state and the long-range Coulombic repulsion of the ionic state. The model achieved a remarkable mean absolute error of just 4.57 a.u. for molecular polarizability, highlighting its robust and generalizable learned representation of polarization physics. Finally, we demonstrated our foundation model’s capacity to achieve ab initio accuracy for specific systems through efficient finetuning.
We acknowledge that our model aims to achieve a balance between computational efficiency and accuracy. Several advanced machine learning approaches for long-range interactions may achieve higher accuracy for specific, complex systems, including higher-order charge expansion methods and simultaneous optimization of QEq parameters 68, 69 in each step. However, more accurate methods incur higher computational complexity. The method that dynamically solve charge equilibration parameters at each step require additional cost for MD simulations. It is also observed that the charge equilibration parameters do not change significantly during MD simulations for that method 69, suggesting that expanding polarization charges to higher orders while using fixed, a priori parameters may achieve higher accuracy with tolerable computational overheads. Additionally, several approaches that do not require charge equilibration process provide valuable insights. Latent Ewald Summation 20 captures long-range interactions by learning hidden variables from local descriptors and applying Ewald summation to these variable. LSR-MP 70 employs a fragmentation-based approach with hierarchical message passing between atoms and fragments to learn the long-range interactions. NeuralP3M 71 introduces mesh points alongside atoms to discretize space and capture long-range effects. SO3krates 72 utilizes spherical harmonic coordinates and equivariant attention for interactions at arbitrary length scales. In contrast to these approaches, our foundation model explicitly simulates charge redistribution through a physics-driven framework that maintains a transparent connection to the underlying physical principles of charge polarization and electrostatic interactions. This design choice enhances interpretability while preserving computational efficiency. A promising direction for future work lies in exploring strategic integrations of our framework with these complementary methodologies to develop even more comprehensive and powerful foundation models.
Methods
Model training
To train the foundation equivariance neural network potential, we used the MPtrj Dataset16 sourced from Materials Projects36 as the training dataset. All configurations were calculated using DFT with the PBE73/PBE+U74 exchange-correlation functional and pseudopotential basis. The PBE/PBE+U mixing compatibility correction75 was applied to ensure energy consistency. Due to transferability issues with the Yb element 36, all data containing Yb in the MPtrj dataset were excluded. The higher-order energy terms and the electrostatic Coulomb interaction energies due to charge fluctuations should be deducted from the total potential energy. Appropriately setting the PQEq parameters is crucial for training a universal framework. In particular, the Li parameters were refitted to reproduce the electrostatic interaction energies observed in QM. Detailed information on the fitting process and the PQEq parameters used in this work can be found in the Supplementary Information.
For equivariant neural network training, the NequIP12 model was utilized, incorporating six equivariant layers with node features set at 256 × 0e + 256 × 1e. Spherical harmonic basis functions were used to represent interatomic directions, denoted as 1 × 0e + 1 × 1o. Interatomic distances were described using eight Bessel basis functions, and a cut-off value of 5.0 Å was chosen for building the short-range neighbor list. The dataset is divided into training set, validation set, and test set in a ratio of 8:1:1.
The total loss function ($\mathcal{L}$) of prediction values $\bm{x}^{\rm pred}$ and reference first principles values $\bm{x}^{\rm ref}$ can be expressed as,
$L(\bm{x}^{\rm pred},\bm{x}^{\rm pred})=\sum_{i=1}^{N}\frac{1}{N}(\bm{x}^{\rm pred}_{i}-\bm{x}^{\rm ref}_{i})^{2}$ (2)
$\begin{split}\mathcal{L}&=\omega_{E}\cdot L(\bm{E}^{\rm pred},\bm{E}^{\rm ref})+\omega_{F}\cdot L(\bm{F}^{\rm pred},\bm{F}^{\rm ref})\\ &+\omega_{S}\cdot L(\bm{S}^{\rm pred},\bm{S}^{\rm ref}),\end{split}$ (3)
where energies $\bm{E}$, forces $\bm{F}$ and stresses $\bm{S}$ weights were set as $\omega_{E}=1$, $\omega_{F}=10$, and $\omega_{S}=10$ with quantity units of eV, eV/Å, and eV/Å3. The batch size was selected as 128, and the Adam optimizer was employed with an initial learning rate of 10-3. Trainable parameters were initialized with the random seed 3407. The implementation of the entire neural networks and long-range interactions was based on JAX 0.4.2076. The training was conducted on a single NVIDIA H100-PCIe GPU, utilizing CUDA version 12.3 and Driver version 545.23.06.
Molecular dynamics simulations
The Atomic Simulation Environment (ASE)77 was utilized as the interface for geometry relaxations and MD simulations. Structure relaxations were conducted using the limited-memory BFGS78 method. Both atomic coordinates and cell vectors were optimized concurrently until the forces fell below the convergence threshold of 0.05 eV/Å.
The c-LLZO with a space group of $\rm Ia\bar{3}d$ originated from experimental structures detailed in reference42. The structure was expanded to a 2×2×2 supercell. Subsequently, 180 lithium atoms were randomly placed on the 24(d) sites, and an additional 268 lithium atoms on the 96(h) sites. After relaxation, the structure served as the starting point for MD simulations. The isothermal isobaric (NpT) ensemble simulations were performed with the temperature control via Nosé-Hoover thermostat79, 80 and pressure maintained at 1 atm using Parrinello-Rahman barostat81, 82 to determine the lattice constants at various temperatures. Then, the canonical (NVT) ensemble simulations with the Nosé-Hoover thermostat were utilized to ascertain the lithium diffusion properties within c-LLZO. A time step of 2 fs was used for all NpT and NVT simulations. In the case of NVT simulations, following a 100 ps equilibration period, 2-ns trajectories’ mean squared displacements of Li ions were employed to calculate the self-diffusion coefficients at various target temperatures. Additionally, we conducted an uncertainty analysis of the diffusion coefficients, applying the empirical error estimation method as detailed in reference41.
To construct the phase diagram of the $\rm BaTiO_{3}$ crystal structure, a supercell consisting of a 10×10×10 $\rm R3m$ $\rm BaTiO_{3}$ lattice with 5000 atoms was created. The crystal structure data for $\rm BaTiO_{3}$ was sourced from the Materials Project36. NpT simulations with the Nosé-Hoover thermostat and Parrinello-Rahman barostat were performed with a 1-fs timestep. Throughout the simulations, the pressure was consistently maintained at 1 atmosphere. To regulate temperature, the procedure began at 5 K and involved incrementing the temperature by 10 K every 20 ps. For the comprehensive calculation method of the local polarization, please refer to the Supplementary Information.
We utilized a 13760-atom structure consisting of a ca. 13.6-nm thick $\rm Im\bar{3}m$ lithium metal anode with [100] facet, paired with a ca. 19.1 nm thick $\rm Pnma$ $\beta-\rm{Li_{3}PS_{4}}$ electrolyte with [010] facet to illustrate the SEI formation across realistic spatial and temporal scales. The system was structured with periodic lateral dimensions measuring 3.1 nm by 2.6 nm. The unit cell of anode and electrolyte structures originated from the Materials Project36. The entire cell was subject to periodic boundary conditions, with a 1 nm atomic layer fixed on each side along the x-direction to ensure a singular interface reaction. NpT MD simulations were conducted at a standard temperature of 300 K and a pressure of 1 atm by Nosé-Hoover thermostat and Parrinello-Rahman barostat with a timestep of 2 fs. Initial velocities for the particles were assigned randomly following the Maxwell-Boltzmann distribution, and the simulations were carried out for a duration of 4 ns.
Data and software availability
The dataset is accessible through the Materials Project36, 16. Detailed PQEq parameters are provided in the Supplementary Information. The JAX implementation code is available throught https://github.com/reaxnet/reaxnet.
Acknowledgments
We gratefully acknowledge the financial support from the RGC General Research Fund (Grant No. 17309620), Hong Kong Quantum AI Lab Limited, Air @ InnoHK of Hong Kong Government, Seed Fund for Basic Research for New Staff 2022/23 (University Research Committee of the University of Hong Kong, No. 2201101550), the Guangdong Shenzhen Joint Key Fund (Grant No. 2019B1515120045) and the National Natural Science Foundation of China (Grant No. 22073007).